OpenStudy (anonymous):

Solve the following equation for x. 13x-19y=29

TheSmartOne (thesmartone):

Add 19y on both sides Divide by 13 on both sides

OpenStudy (anonymous):

isolate the x on one side of the equation

OpenStudy (anonymous):

so what would x equal to now?

OpenStudy (anonymous):

3.6923076923

OpenStudy (anonymous):

nooooooooooooo....... did they give you a number for y?

OpenStudy (anonymous):

because it doesnt really look right............. lol

OpenStudy (anonymous):

yes 19 i added 19 to both sides and then divided by 13 and got 3.69230

OpenStudy (anonymous):

@icecreamman

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

well it would look like \[x=\frac{ 19y+29 }{13 }\]

TheSmartOne (thesmartone):

First add 19y to both sides: \(\sf 13x-19y+19y=29+19y\)

OpenStudy (anonymous):

why? @TheSmartOne

OpenStudy (anonymous):

then 19y would cancel out on the left side of the equation. then divide 13 on both side of the equation to isolate x @TheSmartOne

OpenStudy (anonymous):

but when you divide it says 3.6....

OpenStudy (anonymous):

what kind of calculator are you using? @RaeBeach9296

TheSmartOne (thesmartone):

You are forgetting that there are variables...

OpenStudy (anonymous):

yes you would leave x by itself. @RaeBeach9296 never said anything was equal to y

OpenStudy (anonymous):

just a regular calculator i added 29 +19 = 48 divided by 13 = 3.6923076923

OpenStudy (anonymous):

oh!!!!!!!!!!!!!! nonononononono lolololol you dont add 29 and 19 together because the 19 has a y attached to it.

OpenStudy (anonymous):

ohhhhh so what should i divide

OpenStudy (anonymous):

nothing you just leave it as is like the equation i showed u way above

OpenStudy (anonymous):

x =

OpenStudy (anonymous):

x= \[\frac{ 19y +29 }{ 13} is the answer?\]

OpenStudy (anonymous):

yup